Chiaki is organizing a candy hunt for the children in her neighbourhood. She spent $102 to buy 500 large candies and 400 small candies. The ration of the price of a large candy to the price of a small candy is 7:4. Find the prices of one large and one small candy
Answer by mananth(16946)

4 x -7 y = 0 .............1
Total value
500 x + 400 y = 10200 .............2
Eliminate y
multiply (1)by 400
Multiply (2) by 7
1600 x -2800 y = 0
3500 x + 2800 y = 71400
Add the two equations
5100 x = 71400
/ 5100
x = 14
plug value of x in (1)
4 x -7 y = 0
56 -7 y = 0
-7 y = 0 -56
-7 y = -56
y = 8
x= 14 cost of large candies in cents
y= 8 cost of small candies in cents
m.ananth@hotmail.ca